OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!! On what intervals within [0, 2π) is f increasing? (Enter your answer using interval notation. If an answer does not exist, enter DNE.) y=(-√2) cos(πx/2-π/4)

OpenStudy (anonymous):

lets to this the boneheaded way. cosine is increasing on \((-\pi,\pi)\) so \(-\cos(x)\) is increasing on \((0,\pi)\)

OpenStudy (anonymous):

set \[\frac{\pi x}{2}-\frac{\pi}{4}=0\] and solve for x

OpenStudy (anonymous):

you get \[\frac{\pi x}{2}=\frac{\pi}{4}\] \[x=\frac{1}{2}\]

OpenStudy (anonymous):

then set \[\frac{\pi x}{2}-\frac{\pi}{4}=\pi\] \[\frac{\pi x}{2}=\frac{5\pi}{4}\] \[x=\frac{5\pi}{2}\]

OpenStudy (anonymous):

sorry last one should just be \[x=\frac{5}{2}\]

OpenStudy (anonymous):

so increasing on \((\frac{1}{2},\frac{5}{2})\)

OpenStudy (anonymous):

it says it's wrong((

OpenStudy (anonymous):

y=(-√2) cos(πx/2-π/4) y'=+pie/root(2) [sin(πx/2-π/4)] @x=1/2 => y'=0 therefore , f increases from [0,1/2]

OpenStudy (anonymous):

either wrong(

OpenStudy (anonymous):

looks right to me http://www.wolframalpha.com/input/?i=y%3D%28-%E2%88%9A2%29+cos%28%CF%80x%2F2-%CF%80%2F4%29 i am sticking with my answer